OpenStudy (anonymous):

100 plus the product of a number and -2 equals 50. What is the number?

OpenStudy (anonymous):

100 + (number × -2) = 50

OpenStudy (anonymous):

i hope you can go from there

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i got 24

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

\[100-2\times 24=50\]??

OpenStudy (anonymous):

OH LOL WAIT!!!

OpenStudy (anonymous):

\[100+(-2n)=50\]\[100-2n=50\]\[-2n=-50\]\[n=25\]

OpenStudy (anonymous):

24, 25, whatever

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

embarrassing

OpenStudy (anonymous):

nah not really

OpenStudy (anonymous):

25!